If i=radical -1 what is the value of i^223
Answer by stanbon(75887)

If i=radical -1 what is the value of i^223
------------
Dividing 223 by 4 you get 4*55 + 3
------------------------------------
i^223 = (i^4)^55 * i^3
--------
But i^4=1, so you get:
--------
i^223 = 1^55*i^3 = 1*i^3 = -i
================================
Cheers,
Stan H.